OpenStudy (anonymous):

A ketone have an absorption peak centred around 1710cm^-1 . From this info deduce a value for the force constant,k, of the C=O double bond.

OpenStudy (aaronq):

\(E_n=hc\bar\nu=\hbar \sqrt{\dfrac{k}{μ}}(n + 1 )\) \(\mu\)= is the reduced mass

OpenStudy (anonymous):

Dont really understand

OpenStudy (anonymous):

I have 2 equations One with c which is speed of light One without

OpenStudy (aaronq):

find the energy first (with the wavenumber) then find k, the spring constant

OpenStudy (anonymous):

how do i calculate energy?

OpenStudy (anonymous):

i got the reduced mass is 6.857

OpenStudy (anonymous):

Wavelength = frequency x planck constant?

OpenStudy (anonymous):

Energy not wavelength

OpenStudy (anonymous):

From wavenumber convert to wavelength. Using wavelength x frequency = c then use frequency x planck's constant got me the value 3.399x10^-20

OpenStudy (anonymous):